OpenStudy (anonymous):

Find the equation of the line. Write the equation using function notation. Through (9,8); parallel to f(x)=2x-9

OpenStudy (asnaseer):

if the line is parallel to f(x)=2x-9 then what would its sope be?

OpenStudy (asnaseer):

*slope

OpenStudy (anonymous):

f(x) = 2x - 10

OpenStudy (anonymous):

f(x) = 2x + b if it passes through (9, 8). 8 = 18 + b/ b = -10.

OpenStudy (anonymous):

What if it was through, (4,2); perpendicular to f(x)=-4x+4

OpenStudy (anonymous):

f(x) = 1/4x + 18